OpenStudy (anonymous):

In calculus when taking the integral of a constant while using u-sub does it become x or u?

OpenStudy (anonymous):

well it's constant * u, but you have to substitute back in whatever you set to u in terms of x

OpenStudy (anonymous):

so the integral of 1+u would be u+u^2/2?

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

yup ! \(\large \int c ~dx = cx + C\) \(\large \int c ~du = cu + C\) \(\large \int c ~dt = ct + C\) \(\large \int c ~dw = cw + C\)

ganeshie8 (ganeshie8):

dont forget the integration constant^

OpenStudy (anonymous):

thank you the actual problem is a definite integral so i dont have to worry about that

ganeshie8 (ganeshie8):

ohk... cool :)

OpenStudy (anonymous):

If any of you could help me double check my answer of the integral from 0 to pi/2 of cos^5 c dx it would be helpful.

OpenStudy (whpalmer4):

What do you get for \[\int_0^{\frac{\pi}{2}}\cos^5 x\,dx\]

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/2} (1-\sin^2 x)^2 \cos x dx\]

OpenStudy (anonymous):

then i used u-sub for sinx

OpenStudy (whpalmer4):

Wouldn't it be easier to just start out with \(u=\cos x\)?

OpenStudy (anonymous):

If it is i wouldn't know how to do it

OpenStudy (whpalmer4):

oh, maybe not...

OpenStudy (anonymous):

So after i used u-sub and distributed i got \[\int\limits_{0}^{\pi/2} 1-2u^2+u^4 \]

OpenStudy (whpalmer4):

Now, if you do change of variable, you have to also change the limits of integration, unless you're going to put the variable back before evaluating the definite integral

OpenStudy (anonymous):

yeah put it back into the equation before evaluating

OpenStudy (whpalmer4):

but strictly speaking, you shouldn't be writing it with the old limits... Okay, and if you integrate \(1-2u^2 + u^4\)?

OpenStudy (anonymous):

oh ok yeah so i would have to multiply the old limits by cosx correct?

OpenStudy (whpalmer4):

well, you apply the u substitution equation to the limits to find the new limits

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\]

OpenStudy (whpalmer4):

if you did \(u=\sin x\), then your new limits would be \(u_1 = \sin 0\) and \(u_2 = \sin \pi/2\) yes, \[\int_0^1 1-2u^2+u^4\,du\]

OpenStudy (whpalmer4):

you're getting close :-)

OpenStudy (anonymous):

and what i got for the integral was \[u-2u^3/3+u^5/5 \] and then i substituted sinx for u

OpenStudy (whpalmer4):

okay, but you could just use 1 and 0 and stay with u now... in any case, what is your final answer? I have to go in a minute or two...

OpenStudy (anonymous):

oh really? \[\sin(1)-2/3\sin^3(1)+1/5\sin^5(1)\]

OpenStudy (whpalmer4):

No, \[u-\frac{2}{3}u^3+\frac{1}{5}u^5\]Evaluate that at \(u=1\) and subtract the value at \(u=0\)

OpenStudy (anonymous):

but doesnt the F(0) evaluate to 0?

OpenStudy (whpalmer4):

It does, but what's your point?

OpenStudy (anonymous):

that it isn't needed in the final answer i would only need to evaluate F(1)

OpenStudy (whpalmer4):

Okay...

OpenStudy (anonymous):

So I guess im wondering if \[\sin(1)-2/3\sin^3(1)+1/5\sin(1)^5 \] is correct

OpenStudy (whpalmer4):

No, it's not. Look, either you use the original equation with the original limits, or you use the substituted equation with the substituted limits.

OpenStudy (whpalmer4):

They give the same answer if you do it correctly.

OpenStudy (whpalmer4):

Either you evaluate \[\left.u-\frac{2}{3}u^3+\frac{1}{5}u^5\right|_0^1\] or \[\left.\sin x - \frac{2}{3}\sin^3x+\frac{\sin^5x}5\right|_0^{\frac{\pi}2}\]